OpenStudy (anonymous):

medal and fan Using the ratio of perfect squares method, what is square root of 89 rounded to the nearest hundredth? Do not use your calculator to answer this question. Use the ratio of perfect squares method.

OpenStudy (anonymous):

@welshfella

OpenStudy (welshfella):

what are the 2 perfect squares spanning 89?

OpenStudy (welshfella):

9^2 = 81 and ?

OpenStudy (welshfella):

what is the next perfect square after 81?

OpenStudy (anonymous):

@Thesmarterone

OpenStudy (welshfella):

10^2 = 100

OpenStudy (welshfella):

the square root of 89 must be between 9 and 10

OpenStudy (anonymous):

ooo ok

OpenStudy (welshfella):

- beacuse 89 is between 81 and 100

OpenStudy (welshfella):

now 89 is 8 more that 81 and 100-81 = 19 so by the method sqrt 89 = 8 8/19 or as a decimal that is 8.42

OpenStudy (anonymous):

ok I c now thanks